# Learning Operators through Coefficient Mappings in Fixed Basis Spaces

### v1

@@ 4.4 Elastic plate
Generalization Performance of RFM-C2C vs. RFM-P2C. In this example, we compare the generalization error of RFM-C2C and RFM-P2C methods by examining the effect of varying the number of training samples on the test results. The generalization capabilities of both methods under fixed RFM basis spaces are evaluated using training sets of 1000 samples, 5100 samples, and the complete dataset. As shown in Figure 48, the C2C method consistently achieves superior generalization performance and higher accuracy compared to the P2C method across all three training sets. This advantage becomes increasingly evident as the amount of training data decreases, demonstrating the robustness of the C2C method in data-scarce scenarios. Furthermore, the worst-case relative error from the L-shaped Darcy flow test set, obtained with full training data using the C2C method, is illustrated in Figure 48.
Figure 42: *
@@
Figure 44: Darcy flow in L-shaped domain: Sample grid points for functions.
Figure 45: *
@@
Figure 48: Training curve comparison between the RFM-C2C and RFM-P2C methods in three training datasets of varying scales. All three test datasets are identically sized, each comprising 7000 instances. The number of grid points is set to $n_{1}=736\times 2$ and $n_{2}=450$, and the number of RFM basis functions is $m_{1}=1024$ and $m_{2}=1024$.
Figure 49: *
Figure 50: *
Figure 51: *
@@
Figure 52: Darcy flow in L-shape domain: Worst-case sample from the L-shape Darcy flow dataset with 51000 training data. (a) Pointwise comparison between the prediction and the ground truth. (b) Exact solution. (c) Predicted solution.
@@ 4.6 High Dimension Poisson Equation
In this example, we consider operator learning for high-dimensional functions. In this setting, if the finite element method is employed to construct the basis functions, one inevitably encounters the curse of dimensionality, as the number of basis functions required for an accurate approximation grows exponentially with the spatial dimension. Therefore, in such cases, adopting the RFM-C2C framework provides a more natural and efficient alternative. Also, in such cases, the number of grid points typically grows exponentially with the dimension, making conventional point-to-coefficient (P2C) frameworks impractical. By employing random feature models (RFM) as the basis functions, our approach effectively alleviates this issue and reduces the complexity of the neural operator network. To illustrate the capability of our method in addressing high-dimensional problems, we present the following numerical example on the high-dimensional Poisson equation:
@@
where $\Omega=[-1,1]^{d}$ and $f(\mathbf{x})=\alpha\frac{\pi^{2}}{4}\sum_{i=1}^{d}\sin\left(\frac{\pi}{2}x_{i}\right)$, which admits the exact solution
@@
To construct the training dataset, we uniformly sample $31^{d}$ points from the domain $1,1]^{d}$ and choose $8$ random values of $a\in[-1,1]$. The solutions are then evaluated for different values of $a$. In particular, we designate as Generation 1 (Gen1) the case of $2$ test values with $a\in[-1,1]$ that are not included in the training set. Furthermore, we perform extrapolation tests on $10$ test values of $a\in[1,2]$, which lie outside the training range and are designated as Generation 2 (Gen2). When training the RFM to obtain the coefficients, we employ the stochastic gradient descent (SGD) method. At each training step, we use a batch size of $1e5$, i.e., we randomly select $10^{5}$ points from the total $31^{10}$ grid points to compute the RFM coefficients.
In this example, we adopt the random feature model as the basis functions and use the same basis for both the input and output spaces, i.e., $m_{1}=m_{2}$. We compare two settings with $m_{1}=m_{2}=128$ and $m_{1}=m_{2}=8192$, respectively. The results are summarized in Table 1. From the results in Table 1, we observe that our RFM-based C2C method is able to effectively handle high-dimensional problems while achieving strong generalization. Using a smaller number of basis functions can accelerate training, though at the expense of a slight loss in accuracy. The training and testing curves are presented in Figure 53, while the worst-case sample is illustrated in Figure 57.
@@
10 $1.69\mathrm{e\text{-}02}$ $3.00\mathrm{e\text{-}02}$ 1208.5s $6.48\mathrm{e\text{-}03}$ $1.27\mathrm{e\text{-}02}$ 43403.6s
Table 1: High-dimensional Poisson: Relative $L^{2}$ errors between the predicted and exact solutions $u_{i}(\mathbf{x})$ for different numbers of basis functions. The training time refers to the time required to train the neural operator.
Figure 53: High-Dimensional Poisson: Training and testing curves for different numbers of RFM basis functions.
Figure 54: *
Figure 55: *
Figure 56: *
Figure 57: High-Dimensional Poisson: Worst-case sample from the high-dimensional Poisson dataset. Visualization of $u(\mathbf{y})$ along the cross-section $(y_{3},y_{4})$ with all other coordinates fixed as $y_{i}=0$ for $i\neq 3,4$. (a) Pointwise comparison between the prediction and the ground truth. (b) Exact solution. (c) Predicted solution.
@@ 5 Conclusion
We propose a coefficient-to-coefficient (C2C) operator learning framework constructed on fixed basis spaces for both the input and output functions, termed the Fixed-Basis Coefficient-to-Coefficient Operator Network (FB-C2CNet). Unlike conventional neural operator frameworks that take pointwise samples on spatial grids as inputs, our method employs the coefficients associated with a set of pre-selected fixed basis functions. This representation substantially reduces the input dimensionality and the number of trainable parameters, thereby alleviating the training difficulty. Moreover, by fixing the basis functions a priori, the proposed framework improves both the stability and efficiency of the training process.
In this work, we employ two representative types of basis functions: the Random Feature Model (RFM) and the Finite Element Method (FEM). We investigate how different parameter settings influence the spectral properties of the coefficient space, showing that configurations with a higher effective rank and more balanced coefficient variations tend to facilitate training and enhance generalization. Our theoretical analysis further demonstrates that the accuracy of the proposed FB-C2CNet is ultimately bounded by the approximation capability of the chosen output function space $V_{\mathrm{out}}$.
We conduct extensive numerical experiments on a wide range of problems, including linear and nonlinear operators, regular and complex domains, and one-, two-, and high-dimensional settings. The results demonstrate the robustness and versatility of the proposed framework. Comparative studies show that the C2C formulation is easier to train and generalizes better than the traditional point-to-coefficient (P2C) approach. The proposed FB-C2CNet also exhibits strong resolution generalization, maintaining high accuracy across different discretization levels. Moreover, we extend the framework to multi-input/output operator learning, as demonstrated in the elastic plate and 2D Darcy flow problems, where multiple input and output functions are involved. The vector-valued RFM formulation in this setting achieves superior efficiency and accuracy, further validating the flexibility and effectiveness of the proposed approach.
@@ Acknowledgments
The work was supported by the Project of Hetao Shenzhen-HKUST Innovation Cooperation Zone HZQB-KCZYB-2020083.
@@
Appendix A Experiments Details
Learning rate
Figure 58: Learning Rate Schedule. The learning rate decays from $10^{-3}$ to $10^{-6}$ for training stability, with a gradual decrease applied within every 10,000 epochs.
Parameter setting
In our method, the following parameters are primarily taken into consideration.
Table 2: Key quantities considered in this case.
Quantity Description
@@ Appendix A Experiments Details
Datasets Number of input grid points $n_{\text{in}}$
Number of output grid points $n_{\text{out}}$
Number of input basis functions $m_{\text{in}}$
Basis systems Number of output basis functions $m_{\text{out}}$
@@
$N_{\text{train}}$ $N_{\text{test}}$ $n_{in}$ $n_{out}$ $m_{in}$ $m_{out}$ $R_{m}$ $M_{p}$ cutcut
1D Darcy Flow 1000 800 2000 2000 128 32 3 (3) [8] ([4]) 1e-2
2D Darcy Flow (regular) 2400 600 19881 19881 1024 2048 0.1 (3) [8,8] ([8,8]) 1e-1
2D Poisson Equation 1600 400 1780 1780 256 400 3 (3) [2,2] ([8,8]) 1e-2
@@
2D Darcy Flow (Lshaped) $51000$ 7000 $736\times 2$ 450 1024 1024 3 (3) [16,16] ([8,8]) 1e-2
2D Darcy Flow (Lshaped) $5100$ 7000 $736\times 2$ 450 1024 1024 3 (3) [16,16] ([8,8]) 1e-2
2D Darcy Flow (Lshaped) $1000$ 7000 $736\times 2$ 450 1024 1024 3 (3) [16,16] ([8,8]) 1e-2
@@
The relative accuracy of our method is as followed.
Table 4: Mean Relative Error $L^{2}$ of the testing data.
FEM-C2C RFM-C2C
1D Darcy Flow 4.26e-2 1.20e-2
2D Darcy Flow (Regular) 7.04e-3 6.07e-3
2D Poisson Equation 8.41e-3 6.40e-3
Elastic Plate Equation 6.26e-1 8.92e-3
2D Darcy Flow (Lshaped) 2.94e-1 3.71e-2
10D Poisson Equation - 6.48e-3

### v2

@@ 4.5 Darcy Flow in an L-shaped domain
Generalization Performance of RFM-C2C vs. RFM-P2C. In this example, we compare the generalization error of RFM-C2C and RFM-P2C methods by examining the effect of varying the number of training samples on the test results. We adopt a piecewise exponential learning rate decay: The learning rate for both methods starts from 1e31e-3 and is multiplied by 0.9 every 1000 steps for a total of 10000 training steps. The generalization capabilities of both methods under fixed RFM basis spaces are evaluated using training sets of 1000 samples, 5100 samples, and the complete dataset. As shown in Figure 17, the C2C method consistently achieves superior generalization performance and higher accuracy compared to the P2C method across all three training sets. This advantage becomes increasingly evident as the amount of training data decreases, demonstrating the robustness of the C2C method in data-scarce scenarios. Furthermore, the worst-case relative error from the L-shaped Darcy flow test set, obtained with full training data using the C2C method, is illustrated in Figure 17.
Table 6: Darcy flow in an L-shaped domain: The comparison of vector-valued C2C framework and P2C framework training on different train sets. The training time refers to the time required to train the neural operator.
Training time Test relative $L^{2}$ error Neural network structure
Data number C2C P2C C2C P2C C2C P2C
1000 19.11s 20.59s 253.92e-3 261.67e-3 [512,256,256,1024] [1472,256,256,1024]
5100 20.96s 24.01s 67.04e-3 96.21e-3 [512,256,256,1024] [1472,256,256,1024]
51000 156.48s 186.67s 38.55e-3 42.09e-3 [512,256,256,1024] [1472,256,256,1024]
@@
Figure 16: Darcy flow in an L-shaped domain: Sample grid points for functions.
@@
Figure 17: Darcy flow in an L-shaped domain: Training curve comparison between the RFM-C2C and RFM-P2C methods in three training datasets of varying scales. All three test datasets are identically sized, each comprising 7000 instances. The number of grid points is set to $n_{1}=736\times 2$ and $n_{2}=450$, and the number of RFM basis functions is $m_{1}=512$ and $m_{2}=1024$.
@@
Figure 18: Darcy flow in an L-shaped domain: Worst-case sample from the L-shaped Darcy flow dataset with 51000 training data. (a) Pointwise comparison between the prediction and the ground truth. (b) Exact solution. (c) Predicted solution.
4.6 KdV-Burgers Equation
To illustrate the capacity of the proposed framework for solving nonlinear time-dependent partial differential equations, we investigate the KdV-Burgers equation with periodic boundary conditions:
$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}$ $=\nu\frac{\partial^{2}u}{\partial x^{2}}-\mathcal{D}\frac{\partial^{3}u}{\partial x^{3}},$ $x\in(0,1),\ t\in(0,1),$ (39)
$u(0,t)$ $=u(1,t),$ $t\in[0,1]$
$\frac{\partial u}{\partial x}(0,t)$ $=\frac{\partial u}{\partial x}(1,t),$ $t\in[0,1]$
$u(x,0)$ $=f(x),$ $x\in[0,1]$
where the viscosity coefficient $\nu=0.01$ and the dispersion coefficient $\mathcal{D}=0.001$. The initial condition $f(x)$ is drawn from a Gaussian random field $\mathcal{N}\left(0,25^{2}(-\Delta+25I)^{-4}\right)$, consistent with the boundary constraints. The referenced solutions are generated using a pseudo-spectral approach discretized over a uniform spatial grid comprising 8192 nodes, integrated in time via a forward Euler scheme with a fine temporal resolution, as outlined in [24]. The resolutions of the input function and output function are 101 and $101\times 101$ respectively. The operator maps the initial condition $f(x)$ to the spatiotemporal solution $u(x,t)$:
$G:f(x)\mapsto u(x,t).$
Fully Coupled Spatiotemporal Basis Functions for Time-Dependent Solution Approximation. To effectively capture the spatiotemporal dynamics of the KdV-Burgers equation, we employ non-separable spatiotemporal basis functions according to RFM basis system (9). Specifically, we define the output basis function as
$\phi(t,x)=\sum_{j=1}^{M}a_{j}\sigma(w_{j}t+k_{j}x+b_{j}),$ (40)
where the inner parameters $\{w_{j},k_{j},b_{j}\}_{j=1}^{M}$ are randomly initialized. The initial learning rates are 1e31e-3, 1e41e-4, and 1e51e-5 for three consecutive 10000-step segments, with a decay factor of 0.9 applied every 600 steps within each segment. The mean relative $L^{2}$ error is 3.29e-3 on the test dataset by RFM-C2C, which verifies the effectiveness of the proposed method for time-dependent PDEs. The worst-case scenario for the KdV-Burgers equation is illustrated in Figure 19.
Figure 19: KdV-Burgers equation: Worst-case sample from the KdV-Burgres equation dataset solved by RFM-C2C. (a) Pointwise comparison between the prediction and the ground truth. (b) Exact solution. (c) Predicted solution.
4.7 2D Helmholtz Equation
In contrast to the previous example with strong solutions, the proposed method can learn mappings to weak solution spaces. We consider the Helmholtz equation defined on a 2 dimensional rectangular domain $\Omega=\{(x,y)\in\mathbb{R}^{2}\mid x\in(-1,1),y\in(0,1)\}$ with parameter $\lambda=1$:
$\begin{cases}\Delta u+\lambda u=f(x,y),\quad(x,y)\in\Omega\\ u|_{\partial\Omega}=\alpha\sinh(|x-\beta|)\cos(y^{2}),\end{cases}$ (41)
where $f(x,y)=\alpha\sinh(|x-\beta|)\left((1+\lambda-4y^{2})\cos(y^{2})-2\sin(y^{2})\right)$, $\alpha$ and $\beta$ are two constants. This equation does not admit a strong solution due to the lack of differentiability at $x=0$, though a weak solution exists. The weak solution is given by $u(x,y)=\alpha\sinh(|x-\beta|)\cos(y^{2})$. To construct the dataset, we choose $50$ random values of $\alpha\in U([1,2])$ and $50$ random values of $\beta\in U([-0.5,0.5])$. Thus, the total data number is 2500. We use 2000 samples for training and 500 samples for testing. The resolutions of the input function and output function are both uniformly discretized as $101\times 101$. We learn the operator that maps forcing term $f(x,y)$ to the weak solution $u(x,y)$:
$G:f(x,y)\mapsto u(x,y).$
Weak Solution Operator Learning. Even though solutions may not be smooth, the proposed method can still learn the weak solution operator $G$ by RFM-C2C (FEM-C2C). The output basis functions are constructed as in (9). The number of basis functions is $m_{1}=256$ $(m_{1}=273)$ and $m_{2}=512$ $(m_{2}=558)$. The RFM basis functions are smooth but can approximate non-smooth solutions in the weak sense. The learning rate for both methods starts from 1e31e-3 and is multiplied by 0.9 every 200 steps for a total of 10000 training steps. The main error occurs near the non-differentiable line. To show clearly, we plot the predicted solution and boundary condition comparison on the boundary $y=0$ and $y=1$. The worst-case scenario for the 2D Helmholtz equation is illustrated in Figure 20 (Figure 21). The mean relative $L^{2}$ error is 6.51e-3 (2.13e-3) in the test dataset by RFM-C2C (FEM-C2C), which verifies the effectiveness of the proposed method for weak solution learning.
Figure 20: 2D Helmholtz equation: The worst-case predicted solution obtained by RFM-C2C method. (a) Pointwise comparison between the prediction and the ground truth. (b) Exact solution. (c) Predicted solution. (d) Predicted solution and boundary condition on $y=0$. (e) Predicted solution and boundary condition on $y=1$.
Figure 21: 2D Helmholtz equation: The worst-case predicted solution obtained by FEM-C2C method. (a) Pointwise comparison between the prediction and the ground truth. (b) Exact solution. (c) Predicted solution. (d) Predicted solution and boundary condition on $y=0$. (e) Predicted solution and boundary condition on $y=1$.
@@ 4.8 High Dimension Poisson Equation
In this example, we consider operator learning for high-dimensional functions. In this setting, if the finite element method is employed to construct the basis functions, one inevitably encounters the curse of dimensionality, as the number of basis functions required for an accurate approximation grows exponentially with the spatial dimension. Therefore, in such cases, adopting the RFM-C2C framework provides a more natural and efficient alternative. Also, in such cases, the number of grid points typically grows exponentially with the dimension, making conventional point-to-coefficient (P2C) frameworks impractical. By employing random feature models (RFM) as the basis functions, our approach effectively alleviates this issue and reduces the complexity of the neural operator network. To illustrate the capability of our method in addressing high-dimensional problems, we present the following example on the 10 dimensional ($d=10$) Poisson equation:
@@
where $\Omega=(-1,1)^{d}$ and $f(\mathbf{x})=\alpha\frac{\pi^{2}}{4}\sum_{i=1}^{d}\sin\left(\frac{\pi}{2}x_{i}\right)$, which admits the exact solution
@@
To construct the training dataset, we uniformly sample $31^{10}$ points from the domain $1,1]^{10}$ and choose $8$ random values of $a\in[-1,1]$.
The solutions are then evaluated for different values of $a$. In particular, we designate as Generation 1 (Gen1) the case of $2$ test values with $a\in[-1,1]$ that are not included in the training set. Furthermore, we perform extrapolation tests on $10$ test values of $a\in[1,2]$, which lie outside the training range and are designated as Generation 2 (Gen2). When training the RFM to obtain the coefficients, we employ the stochastic gradient descent (SGD) method. At each training step, we use a batch size of $1e5$, i.e., we randomly select $10^{5}$ points from the total $31^{10}$ grid points to compute the RFM coefficients. While the P2C approach struggles with the immense scale of $31^{10}$, the C2C method effectively addresses this by reducing the input dimension from $31^{10}$ to a manageable size $m_{1}$.
In this example, we adopt the random feature model as the basis functions and use the same basis for both the input and output spaces, i.e., $m_{1}=m_{2}$. We compare two settings with $m_{1}=m_{2}=128$ and $m_{1}=m_{2}=8192$, respectively. The learning rate for both methods starts from 1e31e-3 and is multiplied by 0.9 every 200 steps for a total of 10000 training steps. From the results in Table 7, we observe that our RFM-based C2C method is able to effectively handle high-dimensional problems while achieving strong generalization. Using a smaller number of basis functions can accelerate training, though at the expense of a slight loss in accuracy. The training and testing curves are presented in Figure 22, while the worst-case sample for the Gen1 when $m_{1}=m_{2}=8192$ is illustrated in Figure 23.
Table 7: High-dimensional Poisson equation: Mean relative $L^{2}$ errors between the predicted and exact solutions $u_{i}(\mathbf{x})$ for different numbers of basis functions. The training time refers to the time required to train the neural operator.
@@
10 $1.69\mathrm{e\text{-}02}$ $3.00\mathrm{e\text{-}02}$ 1208.5s $6.47\mathrm{e\text{-}03}$ $1.27\mathrm{e\text{-}02}$ 76663.5s
Figure 22: High-Dimensional Poisson equation: Training curves comparison for different numbers of RFM basis functions.
Figure 23: High-Dimensional Poisson equation: Worst-case sample from the high-dimensional Poisson Gen1 dataset when $m_{1}=m_{2}=8192$. Visualization of $u(\mathbf{y})$ along the cross-section $(y_{3},y_{4})$ with all other coordinates fixed as $y_{i}=0$ for $i\neq 3,4$. (a) Pointwise comparison between the prediction and the ground truth. (b) Exact solution. (c) Predicted solution.
5 Inverse Boundary Problem for the Stokes System
Forward Problem
Let $\Omega_{0}\subset\mathbb{R}^{2}$ be a fixed computational domain and $\mathcal{O}\subset\Omega_{0}$ denote an obstacle, possibly consisting of multiple connected components. The fluid domain is defined by $\Omega=\Omega_{0}\setminus\mathcal{O},\Gamma=\partial\mathcal{O}.$ We consider the steady incompressible Stokes equations in $\Omega$,
$\begin{cases}-\mu\Delta\mathbf{u}+\nabla p=\mathbf{0},&\mathbf{x}\in\Omega,\\ \nabla\cdot\mathbf{u}=0,&\mathbf{x}\in\Omega,\\ \mathbf{u}=\mathbf{0},&\mathbf{x}\in\Gamma\quad\text{(no-slip)},\\ \mathbf{u}=\mathbf{U}_{\infty},&\mathbf{x}\in\partial\Omega_{0}.\end{cases}$ (43)
Here $\mathbf{u}=(u_{x},u_{y})^{\top}$ denotes the velocity field, $p$ is the pressure, $\mu>0$ is the dynamic viscosity, and $\mathbf{U}_{\infty}$ is a prescribed far-field velocity. The obstacle geometry is represented by a signed distance function (SDF) $\phi:\Omega_{0}\to\mathbb{R}$, defined by
$\phi(\mathbf{x})=\begin{cases}-\mathrm{dist}(\mathbf{x},\Gamma),&\mathbf{x}\in\mathcal{O},\\ \phantom{-}\mathrm{dist}(\mathbf{x},\Gamma),&\mathbf{x}\in\Omega,\end{cases}\qquad\Gamma=\{\mathbf{x}\in\Omega_{0}:\phi(\mathbf{x})=0\}.$ (44)
For a given $\phi$, the forward problem (43) admits a unique velocity solution $\mathbf{u}(\mathbf{x};\phi)$. The solution can be obtained by reformulating the problem as a boundary integral equation. Here, we consider a inverse problem setting.
Inverse problem: sparse velocities-to-geometry
In the inverse setting, the full velocity field $\mathbf{u}(\mathbf{x})$ is not available. Instead, only sparse velocity measurements at a limited number of sensor locations $\{\mathbf{x}_{i}\}_{i=1}^{N_{s}}\subset\Omega_{0}$ are observed:
$\mathcal{Y}=\left\{\bigl(\mathbf{x}_{i},\mathbf{u}(\mathbf{x}_{i})\bigr)\right\}_{i=1}^{N_{s}},\qquad\mathbf{u}(\mathbf{x}_{i})=(u_{x}(\mathbf{x}_{i}),u_{y}(\mathbf{x}_{i})).$ (45)
These sensor locations may be confined to a few spatial windows downstream of the obstacle. In the numerical experiments, we consider a two-dimensional computational domain $\Omega_{0}=[-3,3]\times[-3,3].$ The obstacle $\mathcal{O}$ is assumed to be fully contained within the subdomain $1.5,1.5]\times[-1.5,1.5]\subset\Omega_{0}.$ Velocity measurements are available only in several prescribed observation windows located downstream of the obstacle. Specifically, the sensor locations are restricted to the union of the following regions: $1.5,3.0]\times[-3.0,3.0],[0.5,3.0]\times[-3.0,-2.0],[0.5,3.0]\times[2.0,3.0].$ The dataset comprises 2,500 training samples and 500 testing samples, featuring diverse obstacle geometries such as circular, elliptical, and star-shaped shapes, with one to three obstacles of varying sizes and locations.
The inverse problem consists of recovering the obstacle geometry (equivalently, the signed distance function $\phi$) from the sparse velocity observations:
$\text{Given }\mathcal{Y},\quad\text{find }\phi\text{ such that }\mathbf{u}(\mathbf{x}_{i};\phi)\approx\mathbf{u}(\mathbf{x}_{i}),\quad i=1,\dots,N_{s}.$ (46)
The velocity field is encoded using a random feature model (RFM), whereas the obstacle geometry is parameterized by a fixed random radial basis function (RBF) expansion,
$\phi(\mathbf{x})=\sum_{m=1}^{M}c_{m}\,\psi_{m}(\mathbf{x}),\qquad\psi_{m}(\mathbf{x})=\exp\!\left(-\frac{\|\mathbf{x}-\bm{\mu}_{m}\|^{2}}{2\sigma_{m}^{2}}\right),$ (47)
where the centers $\{\bm{\mu}_{m}\}$ and widths $\{\sigma_{m}\}$ are randomly sampled and fixed. The results are shown below:
Figure 24: Inverse boundary problem. Reconstruction results for different obstacle geometries. The red dashed curves denote the ground-truth boundaries, while the black solid curves represent the recovered obstacles. Black dots indicate the sensor locations, with a total of $N_{s}=100$ velocity measurement points. Overall, the relative $L_{2}$ error is 5.562e25.562e-2.
@@ 6 Conclusion
We propose a coefficient-to-coefficient (C2C) operator learning framework built upon fixed input and output basis spaces, termed the Fixed-Basis Coefficient-to-Coefficient Operator Network (FB-C2CNet). In contrast to conventional neural operator frameworks that rely on pointwise evaluations on spatial grids, our framework operates directly on the coefficients associated with a set of pre-selected basis functions. This coefficient-based representation significantly reduces the input dimensionality and the number of trainable parameters, thereby mitigating training difficulty and improving computational efficiency.
Within this framework, we consider several representative choices of basis functions, including Random Feature Model (RFM) bases, Finite Element Method (FEM) bases and radial basis function (RBF). We show that the effective rank of the encoded coefficient matrix plays a crucial role in determining the generalization error of the learned operator. In particular, a larger effective rank of the coefficient representation consistently leads to improved generalization performance. In addition, we provide a theoretical analysis demonstrating that the accuracy of FB-C2CNet is limited by the projection error associated with the prescribed output basis space $V_{\mathrm{out}}$.
We conduct extensive numerical experiments to evaluate the proposed FB-C2CNet across a wide spectrum of operator learning tasks, including linear and nonlinear operators, regular and complex geometries, one-, two-, and high-dimensional problems, as well as inverse settings. The results demonstrate that FB-C2CNet achieves Competitive predictive fidelity with remarkable computational efficiency. Notably, compared to Point-to-Coefficient (P2C) methods, the proposed framework exhibits superior generalization capabilities, particularly in challenging scenarios.
It is noteworthy that by prescribing the basis functions a priori, the proposed framework substantially reduces the number of trainable parameters while maintaining high-fidelity predictive performance. Consequently, FB-C2CNet demonstrates significantly reduced training overhead compared to conventional neural operator models. Furthermore, this fixed-basis formulation establishes a natural interface with classical numerical methods, enabling the framework to leverage established discretization techniques and facilitating the development of hybrid algorithms that combine data-driven efficiency with numerical reliability.
@@ Acknowledgments
The work of Y.X and W.H.Z was supported in part by the National Key Research and Development Program of China (No. 2025YFA1016800) and the Project of Hetao Shenzhen-HKUST Innovation Cooperation Zone HZQB-KCZYB-2020083.
@@
Experiments Details
Unless otherwise stated, the learning rate follows the scheduling scheme illustrated in Figure 25.
Figure 25: Learning Rate Schedule. The initial rates are 1e21e-2, 1e31e-3, and 1e41e-4 for three consecutive 10000-step segments, with a decay factor of 0.9 applied every 200 steps within each segment
In our method, the quantities shown in Table 8 are primarily considered. Some parameters for each dataset and the RFM basis systems are shown in Table 9.
Table 8: Key quantities and notation.
Quantity Notation
@@
Datasets Number of input grid points $n_{1}$
Number of output grid points $n_{2}$
Number of input basis functions $m_{1}$
Basis systems Number of output basis functions $m_{2}$
@@
$N_{\text{train}}$ $N_{\text{test}}$ $n_{1}$ $n_{2}$ $m_{1}$ $m_{2}$ $R_{m}$ $M_{p}$ cutcut
1D Darcy Flow 1000 800 2000 2000 128 128 3 (3) [16] ([16]) 1e-1
1D Darcy Flow 1000 800 100 100 128 128 3 (3) [8] ([4]) 1e-2
2D Darcy Flow (regular) 2400 600 19881 19881 512 2048 0.3 (0.3) [8,8] ([8,8]) 1e-1
2D Poisson Equation 1600 400 1780 1780 512 1600 3 (3) [2,2] ([4,4]) 1e-2
@@
2D Darcy Flow (Lshaped) $51000$ 7000 $736\times 2$ 450 512 1024 1 (3) [8,8] ([8,8]) 1e-2
2D Darcy Flow (Lshaped) $5100$ 7000 $736\times 2$ 450 512 1024 1 (3) [8,8] ([8,8]) 1e-2
2D Darcy Flow (Lshaped) $1000$ 7000 $736\times 2$ 450 512 1024 1 (3) [8,8] ([8,8]) 1e-2
KdV-Burgers Equation 1600 400 101 19881 32 1024 3 (1) 4 ([8,8]) 1e-2
2D Helmholtz Equation 2000 500 19881 19881 256 512 3 (1) [4,4] ([4,4]) 1e-8
@@
Inverse Stokes system 2500 500 $100$ $128^{2}$ $64$ $1024$ 1 ([0.03,2.4]) [2,2] (-) 1e-2
Experiments Results
The performance of our method is listed in Table 10 and Table 11. The dash (-) in Table 10 is used to represent cases where no experiment was carried out. The multiplication sign ($\times$) indicates layer repetition: 512×5 denotes five consecutive layers, each having 512 neurons.
Table 10: Mean relative $L^{2}$ error of the testing data.
Scalar FEM-C2C Scalar RFM-C2C Vector FEM-C2C Vector RFM-C2C
1D Darcy Flow ($n_{1}=n_{2}=2000$) 13.67e-3 3.87e-3 - -
2D Darcy Flow (Regular) 6.92e-3 6.47e-3 - -
2D Poisson Equation 4.13e-3 7.69e-3 - -
Elastic Plate Equation 1.68e-3 8.47e-3 624.43e-3 7.81e-3
2D Darcy Flow (Lshaped) - - - 38.55e-3
KdV-Burgers Equation - 3.29e-3 - -
2D Helmholtz Equation 2.13e-3 6.51e-3 - -
10D Poisson Equation - 6.48e-3 - -
Inverse Stokes system - 5.56e-2 - -
Table 11: Performance comparison for different benchmarks. Performance are measured with relative $L^{2}$ errors ($RL^{2}E$) and mean squared errors (MSEMSE). The structure of the neural networks used in each experiment is detailed in the “Net” column of the table. Net for BasisONet and B2B includes input basis generator, transformation network and output basis generator. Net for DeepONet means the branch network++ trunk network ++ bias term for the dot product result.
2D Darcy Flow Time Params $RL^{2}E$ MSEMSE Net
[…]
FNO2D 31m22s 1,188,353 5.43e-3 5.19e-8 modes = 12, width = 32, epochs=500
[…]
2D Poisson time params $RL^{2}E$ MSEMSE Net
[…]
Elastic Plate time params $RL^{2}E$ MSEMSE Net
[…]
LShaped Darcy time params $RL^{2}E$ MSEMSE Net
[…]
KdV-Burgers time params $RL^{2}E$ MSEMSE Net
[…]
2D Helmholtz time params $RL^{2}E$ MSEMSE Net
[…]